Please help me solve this problem: Subtract and simplify. 10 1/6 - 8 9/14
Answer by FrankM(1040)

10 1/6 - 8 9/14
We need a least common multiple of 6 and 14. 6 is 2 * 3 and 14 is 2 * 7. THE LCM is 2 * 3 * 7 = 42.
10 1/6 = 10 7/42
8 9/14 = 8 27/42
Now, I can't subtract 7/42 - 27/42 so we'll change 10 7/42 to 9 49/42 (see why?)
9 49/42 - 8 27/42 = 1 22/42 = 1 11/21